OpenStudy (jojokiw3):

Can someone explain to me why this certain answer is the answer to this question?

OpenStudy (jojokiw3):

The question is express sec (alpha) in terms of a & b. 

OpenStudy (jojokiw3):

The answer is \[\sec \alpha = \frac{ \sqrt{a^2 - b^2} }{ b }\]

OpenStudy (jojokiw3):

Wait I think I have it, but I'm not sure. \[\sec \alpha = \frac{ hypotenuse }{ adjacent }\] and \[hypotenuse = \sqrt{a^2 + b^2} \] and \[adjacent = b\]

OpenStudy (jojokiw3):

Oh god I'm stupid.

OpenStudy (jojokiw3):

Haha sorry for wasting all of you guys' time. o.o

ganeshie8 (ganeshie8):

you're not stupid, you have figured it out on ur own !